OpenStudy (anonymous):

In 2000, the circulation of a local newspaper was 3,250. In 2001, its circulation was 3,640. In 2002, the circulation was 4,100. a. Find the percent of increase in the newspaper’s circulation from 2000 to 2001 and from 2001 to 2002. b. Which period had a higher percent of increase, 2000 to 2001, or 2001 to 2002? For questions 18, remember to show all of the steps that you use to solve the problem. Thank you! :)

OpenStudy (campbell_st):

percentage increase 2000 to 2001 \[= \frac{(3640 - 3250)}{3250} \times 100\] for 2001 to 2002 use \[=\frac{4100-3640}{3640} \times 100\]